OpenStudy (anonymous):

6/w = 9/w - 1/5, solve

OpenStudy (anonymous):

6/w - 9/w = -1/5 -3/w = -1/5 -15 = -w 15 = w

OpenStudy (anonymous):

Excellent, thanks once again.

OpenStudy (anonymous):

pls click GOOD ANSWER button!!